# ARM-FM: Automated Reward Machines via Foundation Models for Compositional Reinforcement Learning

### v2

@@ 2.1 Language-Aligned Reward Machines
We present a framework to automatically construct LARMs from language-and-image-based task descriptions by iteratively prompting an FM, as is illustrated in Figure 3. More specifically, to progressively refine the RM specification, we employ N rounds of self-improvement using paired generator and critic FMs (Tian et al., 2024). In practice, we find that FM-generated reward machines are both interpretable and easily modifiable, as they follow a natural language specification. Figure 4 illustrates an automatically-constructed LARM for the UnlockToUnlock task, including a text-based description of the RM (left), FM-generated labeling functions $\mathcal{L}$ (middle), and natural RM state instructions and embeddings $l_{u}$ (right). All RMs and labeling functions used in this work are shown in Appendix 6.6 and 6.7.
@@ 2.2 Reinforcement Learning with LARMs
During training, after the agent executes an action $a_{t}\sim\pi(s_{t},z_{u_{t}})$, the underlying MDP transitions to $s_{t+1}$ and returns a reward $R_{t}$. The labeling function $\mathcal{L}(s_{t+1},a_{t})$ determines if a symbolic event has occurred, which may induce a LARM transition $u_{t+1}=\delta(u_{t},\mathcal{L}(s_{t+1},a_{t}))$, as well as an additional reward $R^{\text{RM}}_{t}$. The sum of the MDP and RM rewards are then used for learning: $R^{\text{total}}_{t}=R_{t}+R^{\text{RM}}_{t}$. This complete training procedure, adapted for a DQN agent, is formalized in Appendix 6.3.
@@ 3 Empirical Results
We present a series of experiments designed to evaluate the effectiveness and scalability of our method: we test generalization and long-horizon planning in sparse-reward settings with the MiniGrid and BabyAI suites (Chevalier-Boisvert et al., 2023) (Section 3.1), we evaluate scalability with a resource-gathering task in a 3D, procedurally generated Minecraft world from Craftium (Malagón et al., 2024) (Section 3.2), and we demonstrate the applicability of our approach to create RMs that work in continuous control in challenging robotics tasks from Meta-World (McLean et al., 2025) (Section 3.3). Finally, we use XLand-MiniGrid (Nikulin et al., 2024) to evaluate the generalization capabilities of RL agents trained with LARMs (Section 3.4). Screenshots of these environments are shown in Figure 5 and environment-specific details in Appendix 6.2.
We report results averaged over 3 independent random seeds, with shaded regions and error bars indicating one standard deviation. Comprehensive details for all environments as well as additional results are presented in Appendix 6.2, details on the baselines used and additional ablations in 6.4, and hyperparameters in 6.8.
@@
Figure 5: An overview of the diverse environments used to evaluate ARM-FM. (a) MiniGrid environments test long-horizon planning in sparse-reward settings. (b) Craftium, a 3D world inspired by Minecraft, evaluates how ARM-FM scales with increasing task complexity. (c) Meta-World provides a benchmark for robotic manipulation with continuous control. (d) XLand-MiniGrid is used to evaluate multi-task learning and generalization.
@@ 3.1 Sparse Reward Tasks
We first evaluate our method on the MiniGrid suite of environments, which are challenging due to sparse rewards. For these experiments, we use DQN (Mnih et al., 2013) as the base agent and compare our method (DQN+RM) with a baseline with intrinsic motivation (DQN+ICM) (Pathak et al., 2017), an LLM-as-agent baseline from Paglieri et al. (2024), as well as an unmodified DQN. In Appendix 6.4.4 we include results comparing to a VLM-as-reward-model baseline proposed by Rocamonde et al. (2023). The LARMs used to train our DQN+RM agent are shown in Appendix 6.6.
@@
Our agent successfully solves a suite of challenging exploration tasks where all baselines fail. We first present results on the DoorKey task in Figure 6, showing performance across increasing grid sizes. Our method consistently outperforms the baselines in all settings, including fixed maps (top row) and procedurally generated maps where the layout is randomized each episode (bottom row). We provide an additional analysis of our method’s success in Appendix 6.4.2.
@@ 3.3 Robotic Manipulation
Our framework can automate the complex task of reward engineering for robotic manipulation, providing dense supervision with a FM-generated LARM. We evaluate this capability in continuous control domains using the Meta-World benchmark (McLean et al., 2025), where designing dense reward functions typically requires extensive hand-engineering of low-level signals (e.g., joint angles). Our approach bypasses this difficulty entirely. The resulting reward machine offers richer learning signals than sparse rewards, enabling the agent to make more progress. As demonstrated in Figure 8, our method achieves higher success rates than learning from sparse rewards alone, using SAC (Haarnoja et al., 2018) as the base agent. Additional experiments on Meta-World are provided in Appendix 6.2.3.
@@ 3.4 Generalization through language embeddings
Figure 9: Ablation study on the components of our framework. A single Rainbow agent is trained on an increasing number of simultaneous tasks. While the baseline fails to generalize, and ablations show partial success, only our full method, which uses both LARM rewards and state embeddings, maintains a high success rate as the number of tasks grows.
A key design choice in our framework is to condition the agent’s policy on the language embeddings of the current RM state, which contrasts with prior work that uses separate policies that do not permit knowledge sharing (Alsadat et al., 2025). In the following, we (i) ablate the roles of the LARM rewards and state embeddings in enabling an RL agent to learn robust, multi-task policies; and (ii) demonstrate how the compositional structure of LARMs leads to zero-shot generalization on unseen tasks.
Both structured rewards and language-based state conditioning are essential for learning a robust, multi-task policy. To disentangle the benefits of the LARM’s reward structure from the state embeddings, we conduct an ablation study in a multi-task setting. We train a single Rainbow DQN agent on an increasing number of simultaneous XLand-MiniGrid tasks and measure its average success rate. As shown in Figure 9, the baseline agent fails to generalize as the number of tasks increases. Providing the policy with only the state embeddings gives it a weak learning signal that degrades quickly. Conversely, providing only the LARM rewards enables multi-task learning, but the policy struggles as it is unaware of the active sub-goal. Our full method, which uses both the dense rewards from the LARM and the state embeddings to condition the policy, is robust and maintains high performance even when trained on 10 simultaneous tasks.
@@ 5 Related Work
Reward Machines in Reinforcement Learning. RMs are a formal language representation of reward functions that expose the temporal and logical structure of tasks, thus enabling decomposition, transfer, and improved sample efficiency in learning (Icarte et al., 2018, 2022). For these reasons, RMs and related formal methods for task specification have been applied to address diverse challenges, from multiagent task decomposition (Neary et al., 2021; Smith et al., 2023) to robotic manipulation and task planning (Camacho et al., 2021; He et al., 2015; Cai et al., 2021). Recent work continues to broaden their applicability, by studying extensions that increase their expressivity (Varricchione et al., 2025), and by addressing uncertainty in symbol grounding and labeling functions (Li et al., 2024, 2025). While RMs can be difficult to design for non-experts, Toro Icarte et al. (2019) and Xu et al. (2020) propose methods that simultaneously learn RMs and RL policies, if the RM is unknown a priori. More closely related to our work, Alsadat et al. (2025) recently proposed to use language models to generate RMs for RL training. However, the above works treats RM states as isolated symbols, requiring carefully-defined mappings between states in different RMs to enable policy re-use. By contrast, ARM-FM not only generates RMs from natural language task descriptions, but also introduces a natural mechanism for connecting RM states by embedding their associated subtask descriptions in a shared latent space. Conditioning the policy on these language embeddings can thus enable knowledge transfer across similar subtasks, even when they occur in different RMs.
@@
Foundation Models in Decision-Making. The emergence of FMs has inspired two main lines of work in sequential decision-making. The first uses FMs directly as autonomous agents (Paglieri et al., 2024). Approaches such as ReAct (Yao et al., 2023) and Voyager (Wang et al., 2023) employ large language models (LLMs) to perform reasoning, planning, and acting. While these systems demonstrate strong capabilities in complex domains, they heavily depend on environment abstractions (e.g., textual interfaces or code as actions) that bypass many of the low-level perception and control challenges central to RL.
@@
A second line of research integrates FMs with RL training by using them to provide auxiliary signals such as high-level goals (Ahn et al., 2022) or reward feedback. For example, Motif (Klissarov et al., 2023) elicits trajectory-level preferences from FMs and distills them into a reward model. ONI (Zheng et al., 2024) aggregates asynchronous LLM feedback into a continuously updated reward function. Eureka (Ma et al., 2023) leverages evolutionary strategies to generate programmatic reward functions, which are then used to train downstream policies. ELLM uses pretrained LLMs to suggest plausibly useful goals and trains RL agents with goal-reaching rewards. These approaches illustrate the potential of injecting FM knowledge to shape RL objectives. However, the outputs are typically limited to an opaque reward model, rather than a structured, compositional representation of the task.
@@
Our work differs in the structure of the FM–RL interface. We employ FMs to generate language-aligned RMs: structured, compositional, and interpretable representations of task reward functions. This formulation combines the expressivity of FMs with the explicit, modular decomposition, and human-in-the-loop refinement enabled by RMs, offering a principled path toward hierarchical and interpretable RL. Additionally, our method does not depend on specific environment abstractions (Wang et al., 2023) or the availability of a natural language interface (Klissarov et al., 2023).
Exploration. Sparse rewards present a fundamental exploration challenge in RL, as discovering rewarding states via random search is often intractable in high-dimensional environments (Sutton et al., 1998). Classic strategies based on state-visitation counts or optimism provide theoretical guarantees but scale poorly to complex state spaces where such counts are infeasible (Strehl and Littman, 2008; Auer et al., 2008; Bellemare et al., 2016). A dominant paradigm in deep RL is intrinsic motivation (Yuan et al., 2024), which augments sparse rewards with a dense, task-agnostic signal. These methods are typically either knowledge-based, rewarding an agent for improving its world model’s predictions (Pathak et al., 2017; Schmidhuber, 1991; Sekar et al., 2020), or novelty-based, rewarding visits to unfamiliar states (Bellemare et al., 2016; Burda et al., 2018; Ostrovski et al., 2017). However, their task-agnostic nature can lead to inefficient exploration misaligned with the true objective, a failure mode known as the “noisy TV” problem (Friston et al., 2012). To make exploration more directed, goal-oriented and hierarchical methods reframe the problem as reaching self-generated goals or learning a repertoire of skills (Andrychowicz et al., 2017; Eysenbach et al., 2018; Klissarov and Machado, 2023). While more structured, these typically rely on complex unsupervised objectives to discover useful goals, which can be unstable (Klissarov et al., 2025). In contrast, our work leverages explicit knowledge from natural language to construct a task-aware reward and state structure, avoiding the pitfalls of purely task-agnostic or unsupervised exploration.

### v3

@@ 2.1 Language-Aligned Reward Machines
We present a framework to automatically construct LARMs from language-and-image-based task descriptions by iteratively prompting an FM, as is illustrated in Figure 3. More specifically, to progressively refine the RM specification, we employ N rounds of self-improvement using paired generator and critic FMs (Tian et al., 2024). A human may optionally intervene by approving the output or providing corrective feedback (see Appendix A.4 for details). In practice, we find that FM-generated reward machines are both interpretable and easily modifiable, as they follow a natural language specification. Figure 4 illustrates an automatically-constructed LARM for the UnlockToUnlock task, including a text-based description of the RM (left), FM-generated labeling functions $\mathcal{L}$ (middle), and natural RM state instructions and embeddings $l_{u}$ (right). All RMs and labeling functions used in this work are shown in Appendix A.9 and A.10. While we use code to define labeling functions in this work, the ARM-FM framework is general, supporting any boolean predicate (e.g. formal logic, or queries to other FMs).
@@ 2.2 Reinforcement Learning with LARMs
During training, after the agent executes an action $a_{t}\sim\pi(s_{t},z_{u_{t}})$, the underlying MDP transitions to $s_{t+1}$ and returns a reward $R_{t}$. The labeling function $\mathcal{L}(s_{t+1},a_{t})$ determines if a symbolic event has occurred, which may induce a LARM transition $u_{t+1}=\delta(u_{t},\mathcal{L}(s_{t+1},a_{t}))$, as well as an additional reward $R^{\text{RM}}_{t}$. The sum of the MDP and RM rewards are then used for learning: $R^{\text{total}}_{t}=R_{t}+R^{\text{RM}}_{t}$. This complete training procedure, adapted for a DQN agent, is formalized in Appendix A.3. The effectiveness of well-designed LARMs yields theoretical guarantees (see Appendix A.5), ensuring that the generated reward structure preserves the optimal policy of the original sparse task.
@@ 3 Empirical Results
Figure 5: Evaluation environments. (a) MiniGrid: tests long-horizon planning with sparse rewards. (b) Craftium: scaling complexity in a 3D Minecraft-inspired world. (c) Meta-World: continuous robotic manipulation. (d) XLand-MiniGrid: tests multi-task generalization.
We present a series of experiments designed to evaluate the effectiveness and scalability of our method: we test generalization and long-horizon planning in sparse-reward settings with the MiniGrid and BabyAI suites (Chevalier-Boisvert et al., 2023) (Section 3.1), we evaluate scalability with a resource-gathering task in a 3D, procedurally generated Minecraft world from Craftium (Malagón et al., 2024) (Section 3.2), and we demonstrate the applicability of our approach to create RMs that work in continuous control in challenging robotics tasks from Meta-World (McLean et al., 2025) (Section 3.3). Finally, we use XLand-MiniGrid (Nikulin et al., 2024) to evaluate the generalization capabilities of RL agents trained with LARMs (Section 3.4). Screenshots of these environments are shown in Figure 5 and environment-specific details in Appendix A.2. We used GPT-4o (Hurst et al., 2024) to generate all LARM components for all tasks, with the exception of the 1,000 LARMs for XLand-MiniGrid. These were generated using various open-source FMs of different scales for our ablation study.
We report results averaged over 3 independent random seeds, with shaded regions and error bars indicating one standard deviation. Comprehensive details for all environments as well as additional results are presented in Appendix A.2, details on the baselines used and additional ablations in A.7, and hyperparameters in A.11.
@@ 3.1 Sparse Reward Tasks
We first evaluate our method on the MiniGrid suite of environments, which are challenging due to sparse rewards. For these experiments, we use DQN (Mnih et al., 2013) as the base agent and compare our method (DQN+RM) with a baseline with intrinsic motivation (DQN+ICM) (Pathak et al., 2017), ReAct (Yao et al., 2023) – an LLM-as-agent baseline which generates reasoning traces as it acts in the environment (Paglieri et al., 2024) as well as an unmodified DQN. In Appendix A.7.4 we include results comparing to a VLM-as-reward-model baseline proposed by Rocamonde et al. (2023). The LARMs used to train our DQN+RM agent are shown in Appendix A.9.
@@
Our agent successfully solves a suite of challenging exploration tasks where all baselines fail. We first present results on the DoorKey task in Figure 6, showing performance across increasing grid sizes. Our method consistently outperforms the baselines in all settings, including fixed maps (top row) and procedurally generated maps where the layout is randomized each episode (bottom row). We provide an additional analysis of our method’s success in Appendix A.7.2.
@@ 3.3 Robotic Manipulation
Our framework can automate the complex task of reward engineering for robotic manipulation, providing dense supervision with a FM-generated LARM. We evaluate this capability in continuous control domains using the Meta-World benchmark (McLean et al., 2025), where designing dense reward functions typically requires extensive hand-engineering of low-level signals (e.g., joint angles). Our approach bypasses this difficulty entirely. The resulting reward machine offers richer learning signals than sparse rewards, enabling the agent to make more progress. As demonstrated in Figure 8, our method achieves higher success rates than learning from sparse rewards alone, using SAC (Haarnoja et al., 2018) as the base agent. Additional experiments on Meta-World are provided in Appendix A.2.3.
@@ 3.4 Generalization through language embeddings
A key design choice in our framework is to condition the agent’s policy on the language embeddings of the current RM state, which contrasts with prior work that uses separate policies that do not permit knowledge sharing (Alsadat et al., 2025). In the following, we (i) ablate the roles of the LARM rewards and state embeddings in enabling an RL agent to learn robust, multi-task policies; and (ii) demonstrate how the compositional structure of LARMs leads to zero-shot generalization on unseen tasks. For clarity, we refer here to zero-shot generalization across novel task compositions within the same domain, rather than cross-domain transfer.
Both structured rewards and language-based state conditioning are essential for learning a robust, multi-task policy. To disentangle the benefits of the LARM’s reward structure from the state embeddings, we conduct an ablation study in a multi-task setting. We train a single Rainbow DQN (Hessel et al., 2018) agent on an increasing number of simultaneous XLand-MiniGrid tasks and measure its average success rate. As shown in Figure 9, the baseline agent fails to generalize as the number of tasks increases. Providing the policy with only the state embeddings gives it a weak learning signal that degrades quickly. Conversely, providing only the LARM rewards enables multi-task learning, but the policy struggles as it is unaware of the active sub-goal. Our full method, which uses both the dense rewards from the LARM and the state embeddings to condition the policy, is robust and maintains high performance even when trained on 10 simultaneous tasks.
113355101000.20.40.60.81.0Number of Simultaneous EnvironmentsSuccess RateBaselineNo RM RewardNo EmbeddingsOurs
Figure 9: Ablation study of the components of ARM-FM. A Rainbow agent is trained on an increasing number of tasks. While baselines fail to generalize, only our full method (combining LARM rewards and state embeddings) maintains high success as the number of tasks grows.
@@ 5 Related Work
Reward Machines in Reinforcement Learning. RMs are a formal language representation of reward functions that expose the temporal and logical structure of tasks, thus enabling decomposition, transfer, and improved sample efficiency in learning (Icarte et al., 2018; 2022). For these reasons, RMs and related formal methods for task specification have been applied to address diverse challenges, from multiagent task decomposition (Neary et al., 2021; Smith et al., 2023) to robotic manipulation and task planning (Camacho et al., 2021; He et al., 2015; Cai et al., 2021). Recent work continues to broaden their applicability, by studying extensions that increase their expressivity (Varricchione et al., 2025), and by addressing uncertainty in symbol grounding and labeling functions (Li et al., 2024; 2025). While RMs can be difficult to design for non-experts, Toro Icarte et al. (2019) and Xu et al. (2020) propose methods that simultaneously learn RMs and RL policies, if the RM is unknown a priori.
Recent work also explores FM-driven automata. While some approaches treat RM states as isolated symbols Alsadat et al. (2025), requiring careful state mapping for policy re-use, others use FMs with classic algorithms for automaton discovery Vazquez-Chanlatte et al. (2025), which requires expert demonstrations. Our work differs by generating RMs directly from language descriptions, without behavioral examples. Concurrently, methods like RAD embeddings Yalcinkaya et al. (2024) have been proposed to condition the policy on the automaton’s topology. We take a complementary, language-first approach: our FM generates language-aligned embeddings for the meaning of each state, which our results show effectively grounds the policy.
By contrast, ARM-FM not only generates RMs from natural language task descriptions, but also introduces a natural mechanism for connecting RM states by embedding their associated subtask descriptions in a shared latent space. Conditioning the policy on these language embeddings can thus enable knowledge transfer across similar subtasks, even when they occur in different RMs.
@@
Foundation Models in Decision-Making. The emergence of FMs has inspired two main lines of work in sequential decision-making. The first uses FMs directly as autonomous agents (Paglieri et al., 2024). Approaches such as ReAct, (Yao et al., 2023) Voyager (Wang et al., 2023) and SayCan Ahn et al. (2022) employ large language models (LLMs) to perform reasoning, planning, and acting. While these systems demonstrate strong capabilities in complex domains, they heavily depend on environment abstractions (e.g., textual interfaces or code as actions) that bypass many of the low-level perception and control challenges central to RL. In contrast, we use RMs to structure policies for learning agents, solving complex sparse reward tasks beyond the reach of non-learning, in-context methods like ReAct which additionally require high-level textual interfaces.
@@
A second line of research integrates FMs with RL training by using them to provide auxiliary signals such as high-level goals or reward feedback. For example, Motif (Klissarov et al., 2023) elicits trajectory-level preferences from FMs and distills them into a reward model. ONI (Zheng et al., 2024) aggregates asynchronous LLM feedback into a continuously updated reward function. Eureka (Ma et al., 2023) leverages evolutionary strategies to generate programmatic reward functions, which are then used to train downstream policies. ELLM uses pretrained LLMs to suggest plausibly useful goals and trains RL agents with goal-reaching rewards. These approaches illustrate the potential of injecting FM knowledge to shape RL objectives. However, the outputs are typically limited to an opaque reward model, rather than a structured, compositional representation of the task.
@@
Our work differs in the structure of the FM–RL interface. We employ FMs to generate language-aligned RMs: structured, compositional, and interpretable representations of task reward functions. This formulation combines the expressivity of FMs with the explicit, modular decomposition, and human-in-the-loop refinement enabled by RMs, offering a principled path toward hierarchical and interpretable RL. Additionally, our method does not depend on specific environment abstractions (Wang et al., 2023) or the availability of a natural language interface (Klissarov et al., 2023). We provide a detailed comparison with existing methods in Section A.6 (see Table 4).